OpenStudy (anonymous):

stuck, help? 2(x-3) x+5

Parth (parthkohli):

An equals sign?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 2x - 6 = x + 5 }\) Distributing. \(\Large \color{Black}{\Rightarrow 2x = x + 11 }\) \(\Large \color{Black}{\Rightarrow 2x - x= 11}\) \(\Large \color{Black}{\Rightarrow x = 11 }\)

OpenStudy (anonymous):

well this is how the problem is set up AC= 2(x-3) BD= x+5

Parth (parthkohli):

And are they equal?

OpenStudy (anonymous):

haha idk i suck at math :p

Parth (parthkohli):

Are they opposite sides in a rectangle or square? if not then which kinda polygon ;p

OpenStudy (anonymous):

okay i am doing an online class and thats how it was set up no shape or anything.

Parth (parthkohli):

then most probably they are equal.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Yuuup